Test the indicated claim about population proportion. Use the 5 step process for hypothesis testing.
In a sample of 167 children selected randomly from one town, it is found that 37 of them suffer from asthma. At the 0.05 significance level, test the claim that the proportion of all children in the town who suffer from asthma is 11%.
1) Null and Alternative Hypothesis
2) Calculator work
3) Test statistic and P-Value
4) Conclusion about the null hypothesis
5) Final conclusion that addresses the original claim

Answer by Boreal(15235)

Ho: mu=11%
Ha: mu NE 11%
alpha=0.05 P{reject Ho|Ho true}
test statistic is a z
critical value is |z|>1.96
one sample proportion with xbar=37/167=0.2216
z=(0.2216-0.11)/sqrt(0.11*0.89)/167
=0.1116/0.0242
=4.61
strongly reject Ho and conclude the true proportion is not 0.11.
p<0.0001 (4.08 x 10^(-6), although when it is this small, <.0001 is sufficient